The table below shows the total cost (TC) and marginal cost (MC) for Choco Lovers, a perfectly competitive firm producing different quantities of chocolate gift boxes. The market price of a gift box is $5 per box.



Instructions: Enter your answers as whole numbers. For profit per gift box, round your answers to 2 decimal places.



a. Fill in the marginal revenue (MR) and average revenue (AR) columns.







b. Given a price of $5 per gift box, how many gift boxes should Choco Lovers produce?




25
gift boxes



What will be the profit per gift box?


$

c. Suppose that Choco Lovers raises the price of gift boxes to $7 per gift box. How many gift boxes should Choco Lovers produce now?


gift boxes



What will be the new profit per gift box?
1
Expert's answer
2020-03-30T07:20:44-0400

MC= TCn - TCn-1


TR= TRn - TRn-1


Since change in quantity is at the interval of 5. So, after the calculation of MR and MC. It is going to be divided by 5.


Profit maximizing condition is MR=MC.


As it can be seen in the table that MR and MC are equal at 45 units of output.


profit-maximizing quantity will be 45 units.


Profit-maximizing price is going to be $21.


Profit =TR - TC (at price $21 and quantity of output 45 units.)= 945 - 157.5 = 787.5


Hence profit will be $787.5


profit per unit of production will be equal to = 787.5/45 = 17.5$